OpenStudy (anikate):

log100x simplify. how?

OpenStudy (helder_edwin):

\[\large \log(100x)=\log10^2+\log x=2+\log x \] i guess this is it.

OpenStudy (anonymous):

log(x) + log(100) or log(x) + 2log(100)

OpenStudy (anonymous):

There's probably more alternate forms, but any of the above work.

OpenStudy (helder_edwin):

\[\large \log x+\log10^2=\log x+2\log10=\log x+2\]

OpenStudy (anikate):

wait does log10 in 2log10 cancel out??

OpenStudy (anikate):

@helder_edwin